# AlphaViT: A Flexible Game-Playing AI for Multiple Games and Variable Board Sizes

### v2

@@ Abstract
This paper presents novel game-playing AI agents based on the AlphaZero framework, enhanced with Vision Transformer (ViT): AlphaViT, AlphaViD, and AlphaVDA. These agents are designed to play multiple board games of various sizes using a single network with shared weights, thereby overcoming AlphaZero’s limitation of fixed-board-size constraints. AlphaViT employs only a transformer encoder, whereas AlphaViD and AlphaVDA incorporate both transformer encoders and decoders. In AlphaViD, the decoder processes outputs from the encoder, whereas AlphaVDA uses a learnable embeddings as the decoder input. The additional decoder layers in AlphaViD and AlphaVDA provide flexibility to adapt to various action spaces and board sizes. Experimental results show that the proposed agents, trained on either individual games or multiple games simultaneously, consistently outperform traditional algorithms such as Minimax and Monte Carlo Tree Search and approach the performance of AlphaZero, despite using a single deep neural network (DNN) with shared weights. In particular, AlphaViT shows strong performance across all tested games. Furthermore, fine-tuning the DNN using pre-trained weights from small-board games accelerates convergence and improves performance, particularly in Gomoku. Interestingly, simultaneous training on multiple games yields performance comparable to, or even surpassing, single-game training. These results indicate the potential of transformer-based architectures to develop more flexible and robust game-playing AI agents that excel in multiple games and dynamic environments.
@@ 1 Introduction
Artificial intelligence (AI) has advanced remarkably in recent years, demonstrating its potential across a wide range of applications. One area where AI has excelled is in mastering board games, often outperforming top human players. Notable achievements include AI systems that defeat humans in games, such as checkers, chess (Campbell et al., 2002), and Othello (Buro, 1997). A significant turning point occurred in 2016 when AlphaGo (Silver et al., 2016), an AI designed specifically for the Go game, defeated the world’s top players. After this, AlphaZero (Silver et al., 2018) was introduced, demonstrating its capability to master various board games, including Chess, Shogi, and Go. This further highlights AI’s superhuman skills of AI in this domain.
@@
Despite these successes, many current game-playing AI agents suffer from a fundamental limitation: they are typically designed for only a single specific game and cannot play other games. Even when the game rules remain consistent, these agents cannot handle variations in board size. In contrast, humans can easily switch between different board sizes. For example, beginners in Go often start practicing on smaller boards (e.g., $9\times 9$) before progressing to larger boards (e.g., $19\times 19$). However, AI agents like AlphaZero, which are designed for a single specific game and fixed board size, require substantial reprogramming to accommodate such changes.
@@
For AlphaZero, this limitation arises from its deep neural network (DNN) architecture, which requires a fixed input size. AlphaZero’s DNN consists of residual blocks and multilayer perceptrons (MLPs). These components are designed for a specific input size. The output size of the residual blocks varies with changes in input size, creating inconsistencies with the expected MLP input size. Consequently, AlphaZero cannot function properly even with small variations in board size.
@@
To address this limitation, this study proposes replacing residual blocks in the AlphaZero framework with Vision Transformer (ViT) (Dosovitskiy et al., 2021). ViT is an image-classification DNN based on the transformer architecture. ViT divides an image into patches and infers a class of the image from the patches. A key advantage of ViT is its flexibility in handling images of various sizes. This flexibility enables the AlphaZero framework to adapt to various games and board sizes.
@@
This paper presents game-playing agents capable of handling multiple games and variable board sizes using a single DNN. These agents, named AlphaViT, AlphaViD (AlphaViT with a transformer decoder layer), and AlphaVDA (AlphaViD with learnable action embeddings), are based on the AlphaZero framework. The agents predict the value of a game state and policy using a DNN, while their decisions are made using Monte Carlo Tree Search (MCTS). Our computational experiments show that the proposed agents can be simultaneously trained on three games (Connect4, Gomoku, and Othello) using a single DNN with shared weights. Moreover, the proposed agents outperform traditional algorithms such as Minimax and MCTS across various games while approaching the performance of AlphaZero, whether trained on a single game or simultaneously trained on multiple games.
@@ 2 Related work
Game-playing AI agents have reached superhuman performance levels in traditional board games such as Checkers (Schaeffer et al., 1993), Othello (Buro, 1997, 2003), and Chess (Campbell, 1999; Hsu, 1999; Campbell et al., 2002). In 2016, AlphaGo (Silver et al., 2016), a Go-playing AI, has defeated the world’s top Go players, marking the first superhuman-level performance in Go. AlphaGo relies on supervised learning from a large database of expert human moves and self-play data. Subsequently, AlphaGo Zero (Silver et al., 2017) has defeated AlphaGo, relying only on self-play data. In 2018, Silver et al. (2018) have proposed AlphaZero, which has no restrictions on playable games. AlphaZero has outperformed other superhuman-level AIs in Go, Shogi, and Chess. Interestingly, AlphaZero’s capabilities extend beyond traditional two-player perfect information games, with research exploring its potential in more complex scenarios. For example, Hsueh et al. (2018) have shown AlphaZero’s potential in nondeterministic games. Other extensions include handling continuous action spaces (Moerland et al., 2018) and support for multiplayer games (Petosa and Balch, 2019). However, AlphaZero cannot play multiple games simultaneously and handle games with the same rules but different board sizes using a single DNN with shared weights. This limitation is due to the use of the DNN in AlphaZero for policy and value estimations. AlphaZero’s DNN consists of residual blocks and MLPs. While the residual blocks can process input images of various sizes, the MLPs fix their input sizes. Consequently, the DNN is not sufficiently flexible to accommodate variations in board size.
@@
Researchers have improved the AlphaZero framework to overcome its limitations. Wu (2019) has tackled some of these limitations by improving the efficiency of AlphaZero-like training in Go. His model introduces techniques such as playout cap randomization and policy target pruning, which significantly accelerate self-play learning. Wu has used global pooling layers to standardize varying board sizes to a fixed size, enabling the model to estimate values using MLPs effectively. Furthermore, Wu has exclusively utilized convolutional layers for policy estimation, thereby avoiding the need for MLPs for policy estimation. This design enables seamless handling of varying board sizes while efficiently calculating the policy. This innovation marks a significant step towards creating more flexible AI agents capable of adapting to different board configurations. Similarly, Soemers et al. (2023) have explored transfer learning across various board games, employing fully convolutional networks with global pooling to enable effective transfer between games with different board sizes, shapes, and rules. Their approach has demonstrated the potential of convolutional architectures, enhanced with global pooling, to generalize across various game scenarios.
@@
To address the limitation of the AlphaZero framework, we focus on the transformer architecture and solve the problem by integrating transformer architecture into the AlphaZero framework. The transformer architecture, initially developed for natural language processing (Vaswani et al., 2017), has shown remarkable effectiveness in various domains. The transformer architecture has also been successfully applied to image-processing tasks. Transformer-based models have achieved exceptional performance in various image-related tasks, including image classification (Dosovitskiy et al., 2021), semantic segmentation (Xie et al., 2021), video classification (Li et al., 2022), and video captioning (Zhao et al., 2022). ViT, introduced by Dosovitskiy et al. (Dosovitskiy et al., 2021), is a remarkable example of a transformer-based model for image processing. ViT has achieved state-of-the-art performance in image classification at the time of its introduction. A key feature of ViT is its independence from the input image size (Dosovitskiy et al., 2021). Unlike convolutional neural networks, which require fixed-size inputs, ViT can process images of various sizes by dividing them into fixed-size patches and treating each patch as a token in the transformer architecture. This flexibility makes ViT highly adaptable and efficient when handling different image sizes. By incorporating the ViT architecture into the AlphaZero framework, we extend its capability to handle various games with different board sizes, enhancing its flexibility.
@@ 3 AlphaViT, AlphaViD, and AlphaVDA
AlphaZero’s game-playing capability is limited to games with specific board sizes and rules used during training. This limitation stems from AlphaZero’s deep neural network (DNN), which consists of residual blocks followed by multilayer perceptrons (MLPs) for value and policy computations. The MLPs receive input from the final residual block and assume a fixed input size. Consequently, AlphaZero’s performance is restricted to games with predetermined board size.
@@
To address this limitation, AlphaViT, AlphaViD, and AlphaVDA are developed as game-playing AI agents based on the AlphaZero framework but using Vision Transformer (ViT) architecture. These agents use a combination of a DNN and MCTS (Fig. 1). The DNN receives the board state, and outputs the value estimates and the move probabilities (policies). The MCTS searches a game tree using the estimated value and move probability. By incorporating ViT instead of residual blocks, AlphaViT, AlphaViD, and AlphaVDA can overcome the limitations of AlphaZero and play games with varying board sizes and rules.
An overview of AlphaViT’s DNN architecture is shown in Fig. 2. The DNN in AlphaViT is based on ViT, which has no input-size limitation and can classify images even if the input image size differs from the training image size. This flexibility enables AlphaViT to play games with different board sizes using the same network. In AlphaViT, the game boards are fed into ViT. Initially, these inputs are transformed into patch embeddings through a convolutional layer. Using the convolution layer allows for easy adjustment of patch division parameters such as patch size, stride, and padding. The final outputs of the encoder layer are used to compute the value and the move probabilities.
AlphaViD and AlphaVDA incorporate both transformer encoder and decoder layers for calculating value and move probabilities, as shown in 3. In AlphaViD, the decoder layer receives the input derived from the output of the encoder layer. In contrast, AlphaVDA employs learnable action embeddings as inputs to the decoder layer. The final outputs of the encoder layer are used to compute the value, while the final outputs of the decoder layer are used to compute the move probabilities.
Importantly, AlphaViT, AlphaViD, and AlphaVDA can play any game that AlphaZero can because they employ the same game-playing algorithm as AlphaZero.
@@
Figure 1: AlphaViT, AlphaViD, and AlphaVDA follow the same procedural framework as AlphaZero. They take the board state and a current player as inputs and determine the next move using Monte Carlo Tree Search (MCTS). MCTS explores the game tree using the value and policy probabilities provided by the DNN.
Figure 2: Overview of the AlphaViT model architecture. An input board state is divided into fixed-size patches and linearly embedded by a convolutional layer. Position embeddings are added to the patch embeddings. The resulting sequence of the patch embeddings is appended with a value embedding, a game embedding, and a pass embedding. The sequence is fed into a transformer encoder. Finally, a multilayer perceptron (MLP) estimates the value and move probabilities by taking the outputs from the last layer of the transformer encoder as its input.
Figure 3: Architecture of AlphaViD and AlphaVDA models. An input board state is divided into fixed-size patches and linearly embedded using a convolutional layer. Position embeddings are added to the resulting patch embeddings to preserve spatial information. The sequence of patch embeddings is appended with a value embedding and a static game embedding. This input sequence is fed into a transformer encoder. The estimated value is obtained from a multilayer perceptron (MLP) that takes the output corresponding to the value embedding from the last layer of the transformer encoder. The output embeddings are input to the transformer decoder. For AlphaViD, the input embeddings for the transformer decoder are generated from the output embeddings corresponding to the board state by a fully connected layer. For AlphaVDM, the input embeddings of the transformer decoder are generated from learnable action embeddings. Finally, the move probabilities are produced by another MLP, which takes the output of the last layer of the transformer decoder.
@@ 3.1 Architectures of DNNs
Given a game state $s$, AlphaViT’s DNN predicts the value $v(s)$ and the move probability distribution $\bm{p}(s)$, which includes $p(a|s)$ for each action $a$. In a board game, $s$ represents the board state, and $a$ denotes a move.
The input to the DNN is an $H\times W\times(2T+1)$ image stack ${\mathbf{x}}\in\mathbb{R}^{H\times W\times(2T+1)}$, consisting of $12T+$ 1 binary feature planes of size $H\times W$. Here, $H$ and $W$ are the dimensions of the board, and $T$ is the number of history planes. The first $T$ feature planes represent the occupancies of the first player’s discs, where a feature value of 1 means that a disc occupies the corresponding cell and, and 0 means it does not. The following $T$ feature planes represent the occupancies of the second player’s discs. The final feature plane represents the color of the current player’s disc, where 1 denotes the first player and -1 denotes the second player.
@@
The convolutional layer divides the image stack ${\mathbf{x}}$ into $P\times P$ patches with stride $s$ and padding $p$, and generates the patch embeddings ${\mathbf{z}}_{0}$. The patch size $P$ corresponds to the kernel size of the convolutional layer. The sequence of patch embeddings ${\mathbf{z}}_{0}$ is expressed as
${\mathbf{z}}_{0}=[{\mathbf{x}}^{1}_{p}{\mathbf{E}};...;{\mathbf{x}}^{i}_{p}{\mathbf{E}};...;{\mathbf{x}}^{N_{p}}_{p}{\mathbf{E}}],$ (1)
@@
where ${\mathbf{x}}^{i}_{p}\in\mathbb{R}^{P^{2}(2T+1)}$ is the $i$th flattened 2D patch, and ${\mathbf{E}}$ is a trainable embedding tensor with the shape $(P^{2}(2T+1),D)$. The kernel of the convolutional layer acts as the tensor ${\mathbf{E}}$, which maps each patch to a $D$-dimensional embedding space. Here, $N_{p}$ is the number of embeddings, and $D$ is the embedding size.
@@
To retain position information, learnable 2D position embeddings ${\mathbf{E}}_{\mathrm{pos}}$ are added to the patch embeddings. These positional embeddings are scaled according to the board size and hyperparameters to match the patch embeddings ${\mathbf{x}}_{0}$. The position embeddings are incorporated into the token embeddings ${\mathbf{z}}_{0}$ as follows:
${\mathbf{z}}_{0}={\mathbf{z}}_{0}+E_{\mathrm{pos}}.$ (2)
@@
The output size of the transformer encoder layer is determined by the number of input embeddings. For Gomoku, where the action space is $HW$, AlphaViT requires $N_{p}=HW$ embeddings. To achieve this, we set the patch size $P=2k+1$, where $k$ is a positive integer, stride $s=1$, and padding $\mathrm{pad}=\lfloor P/2\rfloor$. In contrast, for Othello, the action space is $HW+1$ to include the pass move. Using the same parameters as Gomoku ($P=2k+1$, $s=1$, and $\mathrm{pad}=\lfloor P/2\rfloor$), AlphaViT requires one additional embedding for the pass move. To address this, we introduce a learnable pass token $\bm{x}_{\mathrm{pass}}$ and append it to the input embeddings ${\mathbf{z}}_{0}$. To estimate the board value, we prepend a learnable value embedding $\bm{x}_{\mathrm{value}}$. Additionally, to enable AlphaViT to recognize different game types, we incorporate a static game embedding $\bm{x}_{\mathrm{game}}$, represented using one-hot encoding. These embeddings are appended to the input embeddings ${\mathbf{z}}_{0}$, resulting in
${\mathbf{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{N_{p}}_{p}\bm{\mathrm{E}};\bm{x}_{\mathrm{pass}}].$ (3)
@@
Position embeddings are added before appending the pass, value, and game embeddings. This approach ensures that the position embeddings can be scaled independently without being affected by embeddings that do not inherently contain positional information.
@@
The sequence $\bm{z}_{0}$ is fed into the transformer encoder, which consists of $L$ transformer encoder layers. The output of the final encoder layer ${\mathbf{z}}_{L}$ consists of $M\times N+3$ vectors. The first vector $z^{0}_{L}$ derived from the value embedding is processed by the value head implemented as a multilayer perceptron (MLP) denoted as $\text{MLP}_{v}$. This head estimates the value $v$ using the following equation:
$v=\tanh(\mathrm{NLP}_{v}(\mathrm{LN}(z^{0}_{L}))),$ (4)
@@
where LN represents layer normalization. The tanh activation constrains the value within the range $(-1,1)$, representing the winning probability.
The vectors $z^{2}_{L},...,x^{WH+2}_{L}$, derived from the board patches and the pass embedding, are processed by the policy head, which is implemented as another MLP ($\mathrm{NLP}_{p}$). The policy head outputs the move probability $p(m,n)$ for each board position $(m,n),0\leq m\leq W,0\leq n<H$ using the sigmoid function. The policy head outputs a vector $\mathbb{R}^{H\times W+1}$ representing the move probabilities
$p(m,n)=\mathrm{Sigmoid}(\mathrm{MLP}_{p}(\mathrm{LN}(z^{m*H+n+2}_{L})))(m,n).$ (5)
For Othello, $p(W,0)$ corresponds to the probability of a pass action. In contrast, for Gomoku, a pass action is not used. For Connect4, since players can only select from $W$ actions, AlphaViT uses only $p(m,0)$.
@@
To decide on a move, AlphaViT employs Monte Carlo Tree Search (MCTS) with Upper Confidence Bound (UCT), using the value and move probabilities computed by the DNN. The MCTS algorithm used in AlphaViT is identical to that of AlphaZero, as detailed in Appendix 1. The hyperparameters for AlphaViT are listed in Appendix 3.

### v3

@@ Abstract
We present three game-playing agents incorporating Vision Transformers (ViT) into the AlphaZero framework: AlphaViT, AlphaViD (AlphaViT with a transformer decoder), and AlphaVDA (AlphaViD with learnable action embeddings). These agents can play multiple board games of varying sizes using a single neural network with shared weights, thus overcoming AlphaZero’s limitation of fixed board sizes. AlphaViT employs only a transformer encoder, whereas AlphaViD and AlphaVDA incorporate both a transformer encoder and a decoder. In AlphaViD, the decoder processes outputs from the encoder, whereas AlphaVDA uses learnable embeddings as the decoder inputs. The additional decoder in AlphaViD and AlphaVDA provides flexibility to adapt to various action spaces and board sizes. Experimental results show that the proposed agents, trained on either individual games or on multiple games simultaneously, consistently outperform traditional algorithms, such as Minimax and Monte Carlo Tree Search. They approach the performance of AlphaZero despite relying on a single deep neural network (DNN) with shared weights. In particular, AlphaViT performs strongly across all evaluated games. Furthermore, fine-tuning the DNN with weights pre-trained on small board games accelerates convergence and improves performance, particularly in Gomoku. Interestingly, simultaneous training on multiple games yields performance comparable to, or even surpassing, that of single-game training. These results indicate the potential of transformer-based architectures for developing more flexible and robust game-playing AI agents that excel in multiple games and dynamic environments.
@@ 1 Introduction
Artificial intelligence (AI) has advanced remarkably in recent years, demonstrating its potential across a wide range of applications. One area where AI has excelled is mastering board games, often outperforming top human players. Notable achievements include AI agents that have defeated humans in games, such as Checkers, Chess (Campbell et al.,, 2002), and Othello (Buro,, 1997). A significant turning point occurred in 2016 when AlphaGo (Silver et al.,, 2016), an AI designed specifically for the game of Go, defeated the world’s top players. Subsequently, AlphaZero (Silver et al.,, 2018) was introduced, demonstrating its capability to master various board games, including Chess, Shogi, and Go. These achievements further highlight AI’s superhuman skill in this domain.
@@
Despite these successes, many current game-playing AI agents suffer from a fundamental limitation: they are typically designed for only a single specific game and cannot play other games. Even within the same game, these agents cannot handle variations in board size. In contrast, humans can easily switch between different board sizes. For example, beginners in Go often start practicing on smaller boards (e.g., $\times 9$ 9) before progressing to larger boards (e.g., $\times 19$ 19). However, AI agents like AlphaZero, which are designed for a single specific game and fixed board size, require substantial reprogramming to accommodate such changes.
@@
For AlphaZero, this limitation arises from its deep neural network (DNN) architecture, which requires a fixed input size. AlphaZero’s DNN consists of residual blocks and multilayer perceptrons (MLPs) and is designed for a fixed input size. The output size of the residual blocks varies with changes in input size, creating inconsistencies with the expected MLP input size. Consequently, AlphaZero fails even with small changes in board size.
@@
To address this limitation, we propose replacing residual blocks in the AlphaZero framework with Vision Transformer (ViT) (Dosovitskiy et al.,, 2021). ViT is an image-classification DNN based on the transformer architecture. ViT divides an image into patches and infers its class from the patches. A key advantage of ViT is its flexibility in handling various image sizes. This flexibility enables the AlphaZero framework to adapt to various games and board sizes.
@@
This paper presents game-playing agents capable of handling multiple games and variable board sizes using a single DNN. These agents, named AlphaViT, AlphaViD (AlphaViT with a transformer decoder), and AlphaVDA (AlphaViD with learnable action embeddings), are based on the AlphaZero framework. The agents predict the value of a game state and policy using a DNN, and choose moves via Monte Carlo Tree Search (MCTS) (Browne et al.,, 2012; Winands,, 2017). Our computational experiments show that the proposed agents can be trained to play three games (Connect 4, Gomoku, and Othello) simultaneously using a single DNN with shared weights. Moreover, the proposed agents outperform traditional algorithms, such as Minimax and MCTS, across various games, while approaching the performance of AlphaZero, whether trained on a single game or multiple games simultaneously. Our goal is not to surpass the state-of-the-art strength of specialized single-game agents, but rather to demonstrate that a single transformer architecture can flexibly handle multiple games and board sizes.
@@ 2 Related work
Game-playing AI agents have reached superhuman performance levels in traditional board games such as Checkers (Schaeffer et al.,, 1993), Othello (Buro,, 1997, 2003), and Chess (Campbell,, 1999; Hsu,, 1999; Campbell et al.,, 2002). In 2016, AlphaGo (Silver et al.,, 2016), a Go-playing AI, defeated the world’s top Go players, marking the first superhuman-level performance in Go. AlphaGo relied on supervised learning from a large database of expert human moves and self-play data. Subsequently, AlphaGo Zero (Silver et al.,, 2017) defeated AlphaGo, relying only on self-play data. In 2018, Silver et al., (2018) proposed AlphaZero, which has no restrictions on playable games. AlphaZero outperformed other superhuman-level AIs in Go, Shogi, and Chess. Interestingly, AlphaZero’s capabilities extend beyond traditional two-player perfect information games, with research exploring its potential in more complex scenarios. For example, Hsueh et al., (2018) showed AlphaZero’s potential in nondeterministic games. Other extensions include handling continuous action spaces (Moerland et al.,, 2018) and supporting multiplayer games (Petosa and Balch,, 2019). However, AlphaZero cannot play multiple games simultaneously, nor can it handle game variants with the same rules but different board sizes using a single DNN with shared weights. This limitation is due to AlphaZero’s DNN architecture for policy and value estimation. AlphaZero’s DNN consists of residual blocks and MLPs. While the residual blocks can process input images of various sizes, the MLPs require fixed input sizes. Consequently, the DNN is not sufficiently flexible to accommodate variations in board size.
@@
Researchers have improved the AlphaZero framework to overcome its limitations. Wu, (2019) tackled some of these limitations by improving the efficiency of AlphaZero-like training in Go. Wu’s model introduced techniques such as playout cap randomization and policy target pruning, which significantly accelerate self-play learning. Wu used global pooling layers to standardize varying board sizes to a fixed size, enabling the model to estimate values using MLPs effectively. Furthermore, Wu exclusively utilized convolutional layers for policy estimation, thereby eliminating the need for MLPs. This design enables seamless handling of varying board sizes while efficiently calculating the policy. This innovation represents a significant step toward creating more flexible AI agents capable of adapting to different board configurations. Similarly, Soemers et al., (2023) explored transfer learning across various board games, employing fully convolutional networks with global pooling to enable effective transfer between games with different board sizes, shapes, and rules. Their approach has demonstrated the potential of convolutional architectures, enhanced with global pooling, to generalize across various game scenarios.
@@
To address these limitations, we propose integrating the transformer architecture into the AlphaZero framework. The transformer architecture, initially developed for natural language processing (Vaswani et al.,, 2017), has proven remarkably effective across domains, including image-processing tasks. Transformer-based models achieve exceptional performance in various image-related tasks, such as image classification (Dosovitskiy et al.,, 2021), semantic segmentation (Xie et al.,, 2021), video classification (Li et al.,, 2022), and video captioning (Zhao et al.,, 2022). ViT, introduced by Dosovitskiy et al. (Dosovitskiy et al.,, 2021), is a remarkable example of a transformer-based model for image processing. ViT achieved state-of-the-art performance in image classification at the time of its introduction. A key feature of ViT is its independence from the input image size (Dosovitskiy et al.,, 2021). Unlike convolutional neural networks, which require fixed-size inputs, ViT can process images of various sizes. It achieves this by dividing each image into fixed-size patches, each treated as a token in the transformer architecture. This flexibility makes ViT highly adaptable and efficient when handling different image sizes. By incorporating the ViT architecture into the AlphaZero framework, we aim to extend its capabilities to handle various games with different board sizes and enhance its flexibility.
Some researchers have explored the use of transformers in game-playing AI. For example, Czech et al., (2024) proposed a variant of the AlphaZero framework, AlphaVile, using an original network architecture with modified lightweight transformer blocks and an original loss function. AlphaVile achieved better performance than AlphaZero in Chess. Ruoss et al., (2024) reached grandmaster level in Chess with a transformer model trained solely on a dataset, without any self-play or game-tree search. Monroe and Chalmers, (2024) proposed a transformer-based architecture for Chess-playing AI. Using a simple transformer architecture, they trained agents that reached a high level of play. These single-game successes suggest that a transformer architecture can replace convolutions; in this study, we take the next step and show that a transformer can simultaneously master multiple games and board sizes with a single set of weights.
@@ 3 AlphaViT, AlphaViD, and AlphaVDA
AlphaZero’s game-playing capability is limited to games with specific board sizes and rules used during training, as discussed in previous sections. This limitation arises from AlphaZero’s DNN architecture, in which MLPs require fixed input sizes. To address this limitation, we propose AlphaViT, AlphaViD, and AlphaVDA as game-playing AI agents based on the AlphaZero framework but using ViT architecture. These agents use a combination of a DNN and MCTS (Figure 1). The DNN receives the board state and outputs a value estimate and move probabilities (policies). The MCTS searches a game tree using the estimated value and move probabilities. By incorporating ViT instead of residual blocks, AlphaViT, AlphaViD, and AlphaVDA can overcome the limitations of AlphaZero and play games that have variable board sizes and rule sets. While AlphaViT employs only a transformer encoder, AlphaViD and AlphaVDA employ both a transformer encoder and a decoder. Importantly, AlphaViT, AlphaViD, and AlphaVDA can play any game that AlphaZero can because they employ the same game-playing algorithm as AlphaZero.
The training method for AlphaViT, AlphaViD, and AlphaVDA is identical to that of AlphaZero, consisting of three stages: self-play, augmentation, and update. In the self-play stage, the agent generates training data by playing games against itself. The augmentation stage applies data augmentation techniques to the training data. Finally, during the update stage, the DNN weights are updated using the augmented training data. The details of the training method are described in Appendix B.
@@
Figure 1: Overview of the decision process. The agents take a stack of board planes and the current player as input and determine the next move using Monte Carlo Tree Search (MCTS). MCTS explores the game tree using the value and policy probabilities provided by the DNN.
@@ 3.1 Architectures of DNNs
An overview of AlphaViT’s DNN architecture is shown in Figure 2. The DNN in AlphaViT is based on ViT, which has no input-size limitation and can classify images even if the input image size differs from the training image size. This flexibility enables AlphaViT to play games with different board sizes using the same network. In AlphaViT, the game boards are fed into ViT. Initially, these inputs are transformed into patch embeddings through a convolutional layer. Using a convolutional layer allows easy adjustment of the patch-division parameters, such as patch size, stride, and padding. The final output of the encoder is used to compute the value and the move probabilities. AlphaViT employs three special learnable embeddings (tokens) to enable flexibility for different games and board sizes:
  • Value token $\mathbf{x}_{\mathrm{value}}$: is consumed by an MLP value head to predict the state value $v(s)\in(-1,1)$, where $s$ is the game state, 111 indicates a certain win, and 1-1- 1 a certain loss.
  • Game token $\mathbf{x}_{\mathrm{game}}$: encodes which game is currently being played. We use one-hot vectors to represent the game types (e.g., Connect 4, Gomoku, Othello).
  • Pass token $\mathbf{x}_{\mathrm{pass}}$: represents the “pass” action that exists in Othello but not in Gomoku or Connect 4. By appending a dedicated learnable vector after all patch embeddings, we can predict the probability of the pass move when the rules require it.
Figure 2: AlphaViT architecture. The input board is divided into patches by a convolutional layer, projected to embeddings, and combined with value, game, and pass tokens. After adding position embeddings, the sequence is processed by a transformer encoder. Outputs are used by MLP heads to predict the state value and move probabilities.
@@
Given a game state $s$, AlphaViT’s DNN predicts the value $v(s)$ and the move probability vector $\bm{p}(s)$, which includes $p(a\mid s)$ for each action $a$. In a board game, $s$ represents the board state, and $a$ denotes a move.
[…]
@@
The convolutional layer divides the image stack ${\mathbf{x}}$ into $P\times P$ patches with stride $\mathrm{str}$ and padding $\mathrm{pad}$, and generates the patch embeddings $\mathbf{z}_{\mathrm{patch}}$. The patch size $P$ corresponds to the kernel size of the convolutional layer. The sequence of patch embeddings $\mathbf{z}_{\mathrm{patch}}$ is defined in Eq. 1 as follows:
$\mathbf{z}_{\mathrm{patch}}=[{\mathbf{x}}^{0}_{p}{\mathbf{E}};\ldots;{\mathbf{x}}^{i}_{p}{\mathbf{E}};\ldots;{\mathbf{x}}^{N_{p}-1}_{p}{\mathbf{E}}],$ (1)
@@
where $N_{p}$ is the number of patches, ${\mathbf{x}}^{i}_{p}\in\mathbb{R}^{P^{2}(2T+1)}$ is the $i$th flattened 2D patch, and ${\mathbf{E}}$ is a trainable embedding tensor with the shape $(P^{2}(2T+1),D)$. Here, $N_{p}=\bigl{(}\bigl{\lfloor}\frac{H+2\,\mathrm{pad}-P}{\mathrm{str}}\bigr{\rfloor}+1\bigr{)}\times\bigl{(}\bigl{\lfloor}\frac{W+2\,\mathrm{pad}-P}{\mathrm{str}}\bigr{\rfloor}+1\bigr{)}$, where $H$ and $W$ are the board height and width, $P$ is the patch size, $\mathrm{str}$ is the stride, and $\mathrm{pad}$ is the padding used in the convolutional layer. The kernel of the convolutional layer acts as the tensor ${\mathbf{E}}$, which maps each patch to a $D$-dimensional embedding space. Here, $D$ is the embedding size.
@@
To retain position information, learnable 2D position embeddings ${\mathbf{E}}_{\mathrm{pos}}$ are added to the patch embeddings. These positional embeddings are scaled according to the board size and hyperparameters to match the patch embeddings $\mathbf{z}_{\mathrm{patch}}$. The resulting position-aware patch embeddings are given by Eq. 2:
$\mathbf{z}^{\mathrm{pos}}_{\mathrm{patch}}=\mathbf{z}_{\mathrm{patch}}+{\mathbf{E}}_{\mathrm{pos}}.$ (2)
@@
The output size of the transformer encoder is determined by the number of input embeddings. For Gomoku, where the action space is $HW$, AlphaViT requires $HW$ embeddings (i.e., $N_{p}=HW$). To achieve this, we set the patch size $P=2k+1$, where $k$ is a non-negative integer, stride $\mathrm{str}=1$, and padding $\mathrm{pad}=\lfloor P/2\rfloor$. Note that when $k>0$, patches overlap, but the total number of patches remains $HW$. In contrast, for Othello, the action space is $HW+1$ to include the pass move. Using the same parameters as Gomoku ($P=2k+1$, $\mathrm{str}=1$, and $\mathrm{pad}=\lfloor P/2\rfloor$), AlphaViT requires one additional embedding for the pass move. To address this, we introduce a learnable pass embedding ${\mathbf{x}}_{\mathrm{pass}}$ by appending it after all patch embeddings (see Eq. 3). To estimate the board value, we prepend a learnable value embedding ${\mathbf{x}}_{\mathrm{value}}$. Additionally, to enable AlphaViT to recognize different game types, we incorporate a static game embedding ${\mathbf{x}}_{\mathrm{game}}$, represented using one-hot encoding. These embeddings are appended to the embeddings $\mathbf{z}^{\mathrm{pos}}_{\mathrm{patch}}$. As a result, the input embeddings $\mathbf{z}_{0}$ to the transformer encoder are defined in Eq. 3 as follows:
$\mathbf{z}_{0}=[{\mathbf{x}}_{\mathrm{value}};{\mathbf{x}}_{\mathrm{game}};\mathbf{z}^{\mathrm{pos}}_{\mathrm{patch}};{\mathbf{x}}_{\mathrm{pass}}].$ (3)
@@
Position embeddings are added before appending the pass, value, and game embeddings. This approach ensures that the position embeddings can be scaled independently, without being affected by embeddings that do not inherently contain positional information.
@@
The sequence $\mathbf{z}_{0}$ is fed into the transformer encoder, which consists of $L$ transformer encoder layers. The output of the final encoder layer $\mathbf{z}_{L}$ has shape $(HW+3)\times D$. The first vector $\mathbf{z}^{0}_{L}$ derived from the value embedding is processed by the value head implemented as a multilayer perceptron (MLP) denoted as $\text{MLP}_{v}$. This head estimates the value $v$ using Eq. 4:
$v=\tanh(\mathrm{MLP}_{v}(\mathrm{LN}(\mathbf{z}^{0}_{L}))),$ (4)
@@
where LN represents layer normalization. The tanh activation constrains the value within the range (1,1)(-1,1)( - 1 , 1 ), representing the state value, where 111 indicates a certain win and 1-1- 1 a certain loss.
[…]
$p(a_{i}\mid s)=\mathrm{Softmax}(\mathrm{MLP}_{p}(\mathrm{LN}(\mathbf{z}_{p})))_{i},$ (5)
[…]
@@
To decide on a move, AlphaViT employs MCTS with Upper Confidence Bound applied to Trees (UCT), using the value and move probabilities computed by the DNN. The MCTS algorithm used in AlphaViT is identical to that of AlphaZero, as detailed in Appendix A. The hyperparameters for AlphaViT are listed in Appendix C.